OpenStudy (anonymous):

solve for x.... 4(7-x)/3 =x

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Where is the problem there? Ask

OpenStudy (anonymous):

i dont think its 16

OpenStudy (anonymous):

4(7-3)=16

OpenStudy (anonymous):

yes it is if x=3

OpenStudy (anonymous):

all you have to do is plug 3 into x

OpenStudy (anonymous):

oh im sry its 4(7 minus x) all of that over 3. then x = ?

OpenStudy (anonymous):

\[4(7-x) =3x = 4.7 -4x \]

OpenStudy (anonymous):

@Marrioncito huh? you are solving for x.. this is basic algebra

OpenStudy (anonymous):

\[7x = 4.7 \] so \[x = 4\]

OpenStudy (anonymous):

It´s an equation and the asked for solving for X...:S

OpenStudy (anonymous):

I don´t know then what are you asking for. I just solved that equation...